V(s) 50.94g/mol + O2(g) 32.00 g/mol -> V2O3(s) 149.88 g/mol


1
Expert's answer
2021-10-15T12:31:57-0400

50.94× 2 = 101.88

149.88-101.88= 48

48/149.88× 100= 32.1%